Chick-Fil-A bonds currently sells for $1,025. They have a 9 year maturity, an 7% annual coupon, and a par value of $1,000.

What is its current yield?

Group of answer choices

8.78%

8.81%

6.83%

7.80%
Answer #1

annual coupon=1000*7%=70

Hence current yield=Annual coupon/Current value

=70/1,025

=6.83%(Approx).